# Tile and Slide : A New Framework for Scaling NeRF from Local to Global 3D Earth Observation

### v1

@@ Abstract
Neural Radiance Fields (NeRF) have recently emerged as a paradigm for 3D reconstruction from multiview satellite imagery. However, state-of-the-art NeRF methods are typically constrained to small scenes due to the memory footprint during training, which we study in this paper. Previous work on large-scale NeRFs palliate this by dividing the scene into NeRFs. This paper introduces Snake-NeRF, a framework that scales to large scenes. Our out-of-core method eliminates the need to load all images and networks simultaneously, and operates on a single device. We achieve this by dividing the region of interest into NeRFs that 3D tile without overlap. Importantly, we crop the images with overlap to ensure each NeRFs is trained with all the necessary pixels. We introduce a novel $2\times 2$ 3D tile progression strategy and segmented sampler, which together prevent 3D reconstruction errors along the tile edges. Our experiments conclude that large satellite images can effectively be processed with linear time complexity, on a single GPU, and without compromise in quality.
@@
Figure 1: We propose a new framework, to scale up NeRFs without scaling up hardware requirements by an out-of-core strategy. Top row represents: Novel View Synthesis (NVS). Second row: the corresponding depth map with highest part in dark blue and lowest in red. The $2\times 2$ window shows the dynamic adjustment to focus on unexplored regions while maintaining persistent information.
@@ 1 Introduction
  • A $2\times 2$ sliding window mechanism to ensure each ray is used at least once in a training batch with all the necessary NeRFs in memory.
@@ 3 Problem Formulation
Tiling NeRFs require special care to avoid catastrophic forgetting. We perform a simple experiment to illustrate the impact of naïve image cropping on a single NeRF. First, we divide the area into a $2\times 2$ grid following the XY (East/North) directions. We call $R$ the ray set that covers the whole scene, and $R_{i}$ a subset of rays that intersects with a sub-division of the scene, with $i\in\{1,\dots,4\}$, indicating the lower left, upper left, lower right, and upper right quadrants. For this experiment we train the NeRF in four sequential rounds, using samples from each subsets of rays $R_{i}$, one at a time.
@@ 4.1 NeRFs Architecture
A NeRF is a network that predicts the color of a scene viewed from a certain point. This is achieved by generating a 3D ray from each pixel of the rendered scene and sampling points along this 3D ray. A NeRF receives the coordinates of these samples in the $[0;1]^{3}$ space mapped to the corresponding ground 3D tile. We chose the NGP representation [38] in which these coordinates are used to query multi-resolution feature vectors. The feature vectors are provided as input to a local density network, that outputs the density estimation at the given coordinate, and a feature vector which we call the density embedding. This embedding, in turn, forms the input of a global color network which is shared across NeRFs. We take inspiration from NeRF-XL [28], who demonstrate the regularizing effect of a global color network.
@@
Figure 3: Individual non-overlapping tiles are defined relative to their specific UTM partitions, ensuring contiguous spatial segmentation. The partitioning of images is based on their ground coverage, taking into account the minimum and maximum altitudes. The highlighted areas in orange and purple, corresponding to area view by one of each area, indicate regions of the image with overlap.
@@ 4.4 Sliding window
In the overhead scenario, given the low incidence angle of the satellite, each ray can intersect at most 3 tiles. We call the rays intersecting more than one 3D tile shared rays (Figure 4). Consider the setting where a single NeRF is trained at a time. In this scenario, if the shared rays are included as training rays, they will not be able to propagate any gradient onto the neighboring NeRFs that are not loaded in memory. This means that objects outside the bounding box will be explained as “floaters” (density estimation errors) inside the bounding box, as shown in Figure 9.
@@
On the other hand, excluding shared rays altogether results in a low information density along the edges of the 3D tile also incurs density estimation errors. This phenomenon occurs systematically along the edges of the overall region of interest (Figure 9). Therefore, without overlap, loading multiple NeRFs into memory becomes necessary at some point during training to properly learn the edges between their domains.
For this reason and to ensure optimal utilization of GPU memory, regardless of its capacity, we propose a consistent approach of training 4 NeRFs at a time, following a $2\times 2$ window. Importantly, we include rays that fall within any of the 4 NeRFs loaded in memory, and exclude rays that intersect with NeRFs not loaded in memory. See Figure 5 for a simple 1D example.
@@
Near the edge of the $2\times 2$ area, the rays that intersect NeRFs not loaded in memory are excluded. This implies a lower information density in this area. If the area has not yet been learned, this introduces reconstruction errors which will later be corrected when the sliding window advances. If the area has already been learned, we find that the low information density does not cause catastrophic forgetting. Applying the standard NeRF color consistency loss [37] on rays that intersect only the NeRF in memory is sufficient to preserve learned information.
After a fixed number of iterations, we slide the $2\times 2$ window across the scene to cover the entire area, following the ”snake” pattern in Figure 6. This implies that each NeRF is trained 1, 2 or 4 times respectively for corners, edges, and central tiles. This strategy maintains high spatial locality by reducing the distance between grid areas to be processed and minimizing the need to reload recently processed data. This strategy ensures that only the 2 adjacent NeRFs models are loaded or unloaded into memory.
Figure 6: Window sliding path in a $4\times 4$ grid.
A key to the success of our method relies on continuing the training rounds without resetting the optimizer. If the rays that the NeRF has already seen are discarded from the training batch, the NeRF suffers from catastrophic forgetting, illustrated in Section 3. Presenting the NeRF with information it has already received is akin to a recall strategy that is commonly used in Continual Learning [7, 40, 44, 53, 61].
@@ 4.5 Segmented sampler
A key component of our algorithm is to sample along an entire ray which is crossing multiple NeRFs. For each NeRF view by a ray, we identify near/far (resp. N/F) rays as shown in Figure 7.
@@ 4.6 Implementation
We release an open-source highly parallelized implementation of our method, leveraging the computational power of modern GPUs to achieve efficient and scalable performance. Our approach includes several key contributions:
@@
Our implementation is built upon known frameworks : Nerfstudio [47] and NerfAcc [29]. By employing custom CUDA kernels, we achieve significant speed-ups and optimizations. The source code and all related data will be available, ensuring reproducibility and further research.
@@ 5.2 Experimental setup
  • Our scaled approach, Snake-NeRF, described as in Section 4.
@@ Methods
Ours: grid $3\times 3$ 19.30/21.80 0.86/0.93 0.06 19.45/26.14 0.80/0.94 0.03 17.3/23.86 0.81/0.94 0.03 19.20/26.02 0.80/0.93 0.03
Ours: grid $4\times 4$ 19.16/23.22 0.87/0.95 0.06 18.96/24.42 0.77/0.91 0.04 17.9/24.10 0.80/0.95 0.04 19.53/26.09 0.81/0.94 0.02
@@ 5.4 Results and Analysis
The final row demonstrates our proposed scaling method. As shown in Table 1, the NVS results closely resemble the unscaled version, achieving similar PSNR SSIM. The depth map contains no visible error patterns along the edges of the boundaries, indicating that our method effectively learns the continuity between neighboring domains of our NeRFs models. In fact, we observe that the differences between Snake-NeRF and the unscaled version (Figure 10, bottom) are concentrated on the areas where the NeRF is uncertain [18]. Given the simplistic NeRF model employed in this experiment, elements such as shadows, transient objects, and particularly water (visible in the top left of the Middle and Bottom depth maps) present inherent modeling challenges. These difficulties lead to instability across various model initializations. This explain the difference in Table 1, between the Reference case and our Scaled method.
@@ 6 Conclusion
Our experiments show that scaling to large satellite images is possible with a NeRF framework, under 2 conditions.
  • Each ray must be trained with all intersecting NeRFs in memory
  • Each NeRF must be continuously trained with sufficient information density across its entire input domain $(X,Y,Z)$
@@ 7 Limitations and Future Work
To validate our scaling algorithm in a controlled scenario, we utilized an open dataset that covers relatively small areas [26]. In fact, there are no publicly available datasets of multi-view satellite imagery over large areas on which we can release reproducible results. Our hope is that the open-source release of our code will enable private owners of such datasets to benefit from our scaling methodology.
Furthermore, our method does not incorporate the latest advancements in NeRFs for 3D Earth Observation, such as explicit modeling of shadows, transient objects, and albedo (or similar inverse rendering algorithms such as Gaussian Splatting [21]). Future work will explore the integration of these improvements within the context of scaling up.
Additionally, our method considers a constant min and max altitude on each tile. This might be very inefficient in case the scene contains very high objects, in which case we should use different altitudes per tile. This would require a first coarse reconstruction to estimate altitude bounds per tile.
@@
We believe Snake-NeRF also represents a step towards the fusion of satellite and other sources of imagery within a shared NeRF representation.

### v2

@@ Abstract
Neural Radiance Fields (NeRF) have recently emerged as a paradigm for 3D reconstruction from multiview satellite imagery. However, state-of-the-art NeRF methods are typically constrained to small scenes due to the memory footprint during training, which we study in this paper. Previous work on large-scale NeRFs palliate this by dividing the scene into NeRFs. This paper introduces Snake-NeRF, a framework that scales to large scenes. Our out-of-core method eliminates the need to load all images and networks simultaneously, and operates on a single device. We achieve this by dividing the region of interest into NeRFs that 3D tile without overlap. Importantly, we crop the images with overlap to ensure each NeRFs is trained with all the necessary pixels. We introduce a novel $\times 2$ 2 3D tile progression strategy and segmented sampler, which together prevent 3D reconstruction errors along the tile edges. Our experiments conclude that large satellite images can effectively be processed with linear time complexity, on a single GPU, and without compromise in quality.
@@
Figure 1: We propose a new framework, to scale up NeRFs without scaling up hardware requirements by an out-of-core strategy. Top row represents: Novel View Synthesis (NVS). Second row: the corresponding depth map with highest part in dark blue and lowest in red. The $\times 2$ 2 window shows the dynamic adjustment to focus on unexplored regions while maintaining persistent information.
@@ 1 Introduction
  • A $\times 2$ 2 sliding window mechanism to ensure each ray is used at least once in a training batch with all the necessary NeRFs in memory.
@@ 2.1 Neural fields for satellite imagery
While Gaussian Splatting [21] has shown promise in 3D reconstruction in satellite imagery [2], its application to large-scale imagery presents significant challenges. Specifically, the approximation of satellite sensors as affine cameras, which works well for smaller images, introduces precision errors at larger scales. This leads to inaccuracies in mapping pixel locations to ground coordinates, which is crucial for accurate ray casting. This limitation is critical in the context of large-scale satellite imagery, where high precision is essential for accurate 3D reconstruction.
@@ 3 Problem formulation
Tiling NeRFs require special care to avoid catastrophic forgetting. We perform a simple experiment to illustrate the impact of naïve image cropping on a single NeRF. First, we divide the area into a $\times 2$ 2 grid following the XY (East/North) directions. We call $R$ the ray set that covers the whole scene, and $R_{i}$ a subset of rays that intersects with a sub-division of the scene, with $i\in\{1,\dots,4\}$, indicating the lower left, upper left, lower right, and upper right quadrants. For this experiment we train the NeRF in four sequential rounds, using samples from each subsets of rays $R_{i}$, one at a time.
@@ 4.1 NeRFs architecture
A NeRF is a network that predicts the color of a scene viewed from a certain point. This is achieved by generating a 3D ray from each pixel of the rendered scene and sampling points along this 3D ray. A NeRF receives the coordinates of these samples in the $[0;1]^{3}$ space mapped to the corresponding ground 3D tile. We chose the NGP representation [38] in which these coordinates are used to query multi-resolution feature vectors. The feature vectors are provided as input to a local density network that outputs the density estimation at the given coordinate, and a feature vector which we call the density embedding. This embedding, in turn, forms the input of a global color network which is shared across NeRFs. We take inspiration from NeRF-XL [28], which demonstrates the regularizing effect of a global color network.
@@
Figure 3: Individual non-overlapping tiles are defined relative to their specific UTM partitions, ensuring contiguous spatial segmentation. The partitioning of images is based on their ground coverage, taking into account the minimum and maximum altitudes. The highlighted areas in orange and purple indicate regions where the projections of the 3D tiles in the image overlap.
@@ 4.4 Sliding window
In the overhead scenario, given the low incidence angle of the satellite, each ray can intersect at most 3 tiles. We call the rays intersecting more than one 3D tile shared rays (Figure 4). Consider the setting where a single NeRF is trained at a time.
@@
In this scenario, if the shared rays are included as training rays, they will not be able to propagate any gradient onto the neighboring NeRFs that are not loaded in memory. This means that objects outside the bounding box will be explained as “floaters” (density estimation errors) inside the bounding box, as shown in Figure 9. On the other hand, excluding shared rays altogether results in a low information density along the edges of the 3D tile also incurs density estimation errors. This phenomenon occurs systematically along the edges of the overall region of interest (Figure 9). Therefore, without overlap, loading multiple NeRFs into memory becomes necessary at some point during training to properly learn the edges between their domains.
For this reason and to ensure optimal utilization of GPU memory, regardless of its capacity, we propose a consistent approach of training 4 NeRFs at a time, following a $\times 2$ 2 window. Importantly, we include rays that fall within any of the 4 NeRFs loaded in memory, and exclude rays that intersect with NeRFs not loaded in memory. See Figure 5 for a simple 1D example.
@@
Near the edge of the $\times 2$ 2 area, the rays that intersect NeRFs not loaded in memory are excluded. This implies a lower information density in this area.If the area has not yet been learned, this introduces reconstruction errors. However, these excluded rays will be accepted and processed when the sliding window advances over them, thereby correcting any reconstruction errors. If the area has already been learned, we find that the low information density does not cause catastrophic forgetting. Applying the standard NeRF color consistency loss [37] on rays that intersect only the NeRF in memory is sufficient to preserve learned information.
Figure 6: Window sliding path in a $\times 4$ 4 grid.
After a fixed number of iterations, we slide the $\times 2$ 2 window across the scene to cover the entire area, following the ”snake” pattern in Figure 6. This implies that each NeRF is trained 1, 2 or 4 times respectively for corners, edges, and central tiles. This strategy maintains high spatial locality by reducing the distance between grid areas to be processed and minimizing the need to reload recently processed data. This strategy ensures that only the 2 adjacent NeRFs models are loaded or unloaded into memory. A key to the success of our method relies on continuing the training rounds without resetting the optimizer. If the rays that the NeRF has already seen are discarded from the training batch, the NeRF suffers from catastrophic forgetting, illustrated in Section 3. Presenting the NeRF with information it has already received is akin to a recall strategy that is commonly used in Continual Learning [7, 40, 44, 53, 61].
@@ 4.5 Segmented sampler
A key component of our algorithm is to sample along an entire ray which is crossing multiple NeRFs. For each NeRF viewed by a ray, we identify near/far (resp. N/F) rays as shown in Figure 7.
@@ 4.6 Implementation
We release an open-source highly parallelized implementation of our method, leveraging the computational power of modern GPUs to achieve efficient and scalable performance thanks to the following contributions :
@@
Our implementation is built upon known frameworks : Nerfstudio [47] and NerfAcc [29]. By employing custom CUDA kernels, we achieve significant speed-ups and optimizations. The source code and all related data will be available, ensuring reproducibility and further research. The training procedure shares the same loss function and optimization strategy as I-NGP [38].
@@ 5.2 Experimental setup
  • Our scaled approach, Snake-NeRF (Section 4).
@@ Methods
Ours: grid $\times 3$ 3 19.30/21.80 0.86/0.93 0.06 19.45/26.14 0.80/0.94 0.03 17.3/23.86 0.81/0.94 0.03 19.20/26.02 0.80/0.93 0.03
Ours: grid $\times 4$ 4 19.16/23.22 0.87/0.95 0.06 18.96/24.42 0.77/0.91 0.04 17.9/24.10 0.80/0.95 0.04 19.53/26.09 0.81/0.94 0.02
@@ 5.4 Results and analysis
The final row demonstrates our proposed scaling method. As shown in Table 1, the NVS results closely resemble the unscaled version, achieving similar PSNR SSIM. The depth map contains no visible error patterns along the edges of the boundaries, indicating that our method effectively learns the continuity between neighboring domains of our NeRFs models. In fact, we observe that the differences between Snake-NeRF and the unscaled version (Figure 10, bottom) are concentrated on the areas where the NeRF is uncertain [18]. Given the simplistic NeRF model employed in this experiment, elements such as shadows, transient objects, and particularly water (visible in the top left of the Middle and Bottom depth maps) present inherent modeling challenges. These difficulties lead to instability across various model initializations. This explains the difference in Table 1, between the Reference case and our Scaled method.
@@ 6 Conclusion
Our experiments show that scaling to large satellite images is possible with a NeRF framework, as long as:
  • Each ray is trained with all intersecting NeRFs in memory
  • Each NeRF is continuously trained with sufficient information density across its entire input domain $(X,Y,Z)$
@@ 7 Limitations and future work
To validate our scaling algorithm in a controlled scenario, we utilized an open dataset that covers relatively small areas [26]. In fact, there are no publicly available datasets of multi-view satellite imagery over large areas on which we can release reproducible results, preventing further comparisons with multiple sets of data. Our hope is that the open-source release of our code will enable private owners of such datasets to benefit from our scaling methodology.
Additionally, our method does not incorporate the latest advancements in NeRFs for 3D Earth Observation, such as explicit modeling of shadows, transient objects, and albedo (or similar inverse rendering algorithms such as Gaussian Splatting [21]). Future work will explore the integration of these improvements within the context of scaling up.
@@
8 Acknowledgements
This work was performed using HPC resources from CNES Computing Center (DOI 10.24400/263303/CNES_C3). The elevations figures was computed using xDEM package (DOI 10.5281/zenodo.4809697). The authors would like to thank the Johns Hopkins University Applied Physics Laboratory and IARPA for providing the data used in this study, and the IEEE GRSS Image Analysis and Data Fusion Technical Committee for organizing the Data Fusion Contest.